OpenStudy (ilovelife):

Part 1: Show the work to solve |5x – 2| 8 Part 2: Describe the graph of the solution in words.

OpenStudy (helder_edwin):

is it an equation or an inequation??

OpenStudy (helder_edwin):

there's something missing between the absolute value and the 8. is it \(=\) or \(\geq\) or \(\leq\) or \(>\) or \(<\) ???

OpenStudy (ilovelife):

it wont let me put it, but it is the one you put after the =

OpenStudy (helder_edwin):

\[ \large |5x-2|\geq8 \] ??

OpenStudy (helder_edwin):

use this \[ \large |a|\geq b\quad\Leftrightarrow\quad a\geq b\vee a\leq-b \]

OpenStudy (ilovelife):

im even more confused now

OpenStudy (helder_edwin):

\[ \large |5x-2|\geq8 \] becomes \[ \large 5x-2\geq8\qquad\text{or}\qquad 5x-2\leq-8 \]

OpenStudy (helder_edwin):

can u solve these two inequations???

OpenStudy (ilovelife):

i got \[x \ge 2\] and \[x \le \frac{ -6 }{ 5 }\]

OpenStudy (ilovelife):

is that correct ?

jimthompson5910 (jim_thompson5910):

\[x \ge 2\] and \[x \le \frac{ -6 }{ 5 }\] are the correct answers You can combine the two to get the compound inequality \[\Large 2 \le x \le -\frac{6}{5}\]